An ice chest at a beach party contains 12 cans of soda at 4.91 °C. Each can of soda has a mass of 0.35 kg and a specific heat capacity of 3800 J/(kg C°). Someone adds a 7.37-kg watermelon at 27.4 °C to the chest. The specific heat capacity of watermelon is nearly the same as that of water. Ignore the specific heat capacity of the chest and determine the final temperature T of the soda and watermelon in degrees Celsius.
Answer #1

Specific heat of soda = C1 = 3800 J/(kg.oC)

Specific heat of water = Specific heat of watermelon = C2 = 4186 J/(kg.oC)

Number of soda cans = n = 12

Mass of each soda can = m1 = 0.35 kg

Mass of the watermelon = m2 = 7.37 kg

Initial temperature of the soda cans = T1 = 4.91 oC

Initial temperature of the watermelon = T2 = 27.4 oC

Final equilibrium temperature = T3

The heat lost by the soda cans is equal to the heat gained by the watermelon.

nm1C1(T3 - T1) = m2C2(T2 - T3)

(12)(0.35)(3800)(T3 - 4.91) = (7.37)(4186)(27.4 - T3)

15960T3 - 78363.6 = 845312.468 - 30850.82T3

46810.82T3 = 923676.068

T3 = 19.73 oC

Final temperature of the soda and watermelon = 19.73 oC